OpenStudy (anonymous):

What is 8.75% written as a fraction in lowest terms? A.8 3/4 B.7/80 C.8/25 D.17/50

OpenStudy (anonymous):

plz help if u can

OpenStudy (anonymous):

\[\frac{ 8 \frac{3}{4} }{ 100 } = \frac{ \frac{32}{4}+ \frac{3}{4} }{ 100 } = \frac{ \frac{35}{4} }{ 100 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dont now what tht means

OpenStudy (anonymous):

from above: \[\frac{ 35 }{ 400 }\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

If you put it in the lowest terms, you get 7/80